Solve the following linear equations. 
(3t - 2)/4 – (2t + 3)/3 = 2/3 – t 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(3t - 2)/4–(2t + 3)/3 = 2/3–t
⇒((3t–2)/4) × 12–((2t + 3)/3) × 12
⇒(3t - 2) × 3–(2t + 3) × 4 = 2 × 4–12t
⇒9t–6–8t–12=8–12t
⇒9t–6–8t–12=8–12t
⇒t–18=8–12t
⇒t + 12t = 8 + 18
⇒13t = 26
⇒t =2

Thanks